OpenStudy (anonymous):

6 3/4 divided by 7 3/8 = ?

OpenStudy (anonymous):

Keep in mind, dividing by fraction is multiplying by the reciprocal.

OpenStudy (lgbasallote):

27/4/59/8 27/4 x 8/59 27/1 x 2/59 54/59

OpenStudy (anonymous):

well it is easy ummm 6*3/4=18/4=9/2 and 7*3/8=21/8 divide them using mean by mean and extreme by extrem:D then the answer is 72/42=12/7